Adger Corporation is a service company that measures its output based on the number of customers served. The company provided the following fixed and variable cost estimates that it uses for budgeting purposes and the actual results for May as shown below:

Fixed Element
per Month
Variable Element per Customer Served Actual Total
for May
Revenue $ 6,100 $ 223,500
Employee salaries and wages $ 68,000 $ 1,500 $ 126,000
Travel expenses $ 600 $ 20,400
Other expenses $ 47,000 $ 44,300

When preparing its planning budget the company estimated that it would serve 35 customers per month; however, during May the company actually served 40 customers.

a. What amount of revenue would be included in Adger’s flexible budget for May?

b. What amount of employee salaries and wages would be included in Adger’s flexible budget for May?

c. What amount of travel expenses would be included in Adger’s flexible budget for May?

d.What amount of other expenses would be included in Adger’s flexible budget for May?

e.  What net operating income would appear in Adger’s flexible budget for May?
Answer #1
a Revenue in flexible budget 244000 =40*6100
b Employee Salaries and wages in flexible budget 128000 =68000+(1500*40)
c Travel expenses in flexible budget 24000 =40*600
d Other expenses in flexible budget 47000
e Net operating income in flexible budget 45000 =244000-128000-24000-47000